A 6.7 kg block is pushed into a spring that sits o
Answer #1

let

m = 6.7 kg

x = 0.16 m

theta = 20 degrees

a) Apply conservation of energy

initial mechanical energy = final mechanical energy

0.5*k*x^2 + m*g*h1 = 0.5*m*v^2 + m*g*h2

0.5*k*x^2 = 0.5*m*v^2 + m*g*(h2-h1)

0.5*k*x^2 = 0.5*m*v^2 + m*g*x*sin(theta)

k = m*v^2/x^2 + 2*m*g*sin(theta)/x

= 6.7*2.1^2/0.16^2 + 2*6.7*9.8*sin(20)/0.16

= 1435 N/m <<<<<<---------------Answer

b) Again apply conservation of energy

initial mechanical energy = final mechanical energy

0.5*k*x^2 + m*g*h1 = m*g*h2

0.5*k*x^2 = m*g*(h2-h1)

0.5*k*x^2 = m*g*d*sin(theta)

d = 0.5*k*x^2/(m*g*sin(theta))

= 0.5*1435*0.16^2/(6.7*9.8*sin(20))

= 0.818 m <<<<<<---------------Answer